I understand how to calculate odds for an accumulator ... eg 1.2 X 1.4 X 1.7 X 1.3 = 3.71 (or 26.95 %) So I have a 26.95 % chance of winning this acca.
But I tend to bet on much longer shots and I want to know when I am likely to win my next bet. So if I bet on let's say 4 matches with odds such as 21 150 12 65 how do I calculate the chance (in %age) of winning at least once?
in percentage the odds are 21 (4.76%) 150 (0.67%) 12 (8.33%) 65 (1.54%)
It's much easier to work out the chance of losing all the bets.
If individual winning chances are: 4.76%, 0.67%, 8.33% and 1.54% Then the individual losing chances are: 95.24%, 99.33%, 91.67% and 98.46%
So: chance of losing them all is .9524 x .9933 x .9167 x .9846 = .8539, or 85.39% Which means the chance of winning at least one is 100 - 85.39 = 14.61%
It's much easier to work out the chance of losing all the bets.If individual winning chances are: 4.76%, 0.67%, 8.33% and 1.54%Then the individual losing chances are: 95.24%, 99.33%, 91.67% and 98.46%So: chance of losing them all is .9524 x .9933 x .
One more question... What would be the probability of all bets succeeding?
Is it 0.0476 x 0.0067 x 0.0833 x 0.0154 = 0.000000409 (about 1 in 250 million?)
Thanks in advance!
Thanks MR Magoo, that's very clear.One more question...What would be the probability of all bets succeeding?Is it 0.0476 x 0.0067 x 0.0833 x 0.0154 = 0.000000409 (about 1 in 250 million?)Thanks in advance!
Making it too hard on yourself on the last one... just go back to multiplying the numbers together like you did for the first acca!
Your number is right, although you've forgotten it's a percentage so it's 1 in 2.5 million (ish)
Making it too hard on yourself on the last one... just go back to multiplying the numbers together like you did for the first acca!Your number is right, although you've forgotten it's a percentage so it's 1 in 2.5 million (ish)